OpenStudy (anonymous):

I need to figure out the total weight of a box containing 11 cans of weighing 12.5 oz each.

jimthompson5910 (jim_thompson5910):

Multiply 11 by 12.5 to find the total weight

jimthompson5910 (jim_thompson5910):

1 can is 1*12.5 = 12.5 oz 2 cans are 2*12.5 = 25 oz 3 cans are 3*12.5 = 37.5 oz etc etc

OpenStudy (andriod09):

do you still need help?

OpenStudy (anonymous):

yes im trying to figure out postage but the postal office online tells me this exceed thier package limit

OpenStudy (andriod09):

lol?

OpenStudy (andriod09):

bigger box?